OpenStudy (anonymous):

A certain medication is eliminated from the bloodstream at a rate of about 12% per hour. The medication reaches a peak level in the bloodstream of 40 mg. Predict the amount, to the nearest tenth of a milligram, of the medication remaining 2 hours after the peak level and 3 hours after the peak level.

OpenStudy (ivettef365):

do you have choices

OpenStudy (anonymous):

No, you have to write out your response.

OpenStudy (ivettef365):

ok, I would take 12% of 40, that would be the amount it eliminates per hour

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

I got 35.2... lol Probably did that wrong

OpenStudy (ivettef365):

to get 12% of 40, you can multiply 12 x .40

OpenStudy (anonymous):

4.8?

OpenStudy (anonymous):

-30.4 is what I got

OpenStudy (ivettef365):

no wait, it is 12% of 40 = 4.80 40 - 4.80 = 35.20 after one hour now 35.20 x .12 = 4.224 now subtract 35.20 - 4.224 = ??? after 2 hours

OpenStudy (anonymous):

30.976

OpenStudy (ivettef365):

the answer is to the nearest 10nth, so after 2 hours = 30.1 now for 3 hours, multiply 30.976 x .12 and subtract from 30.976

OpenStudy (anonymous):

....What... Why woul... okay lemme try XD

OpenStudy (ivettef365):

the question is after 2 hours then after 3 hours

OpenStudy (anonymous):

I got -27.25

OpenStudy (ivettef365):

30.976 - 3.717 = 27.25 so they want the answer to the nearest 10th that would be 27.3

OpenStudy (anonymous):

okay, thank you!

OpenStudy (ivettef365):

yw :)